### v1

# Explicit unit distance graphs with exponential chromatic number and arbitrary girth

Abstract
In 1975 Erdős initiated the study of the following very natural question. What can be said about the chromatic number of unit distance graphs in $\mathbb{R}^{2}$ that have large girth? Over the years this question and its natural extension to $\mathbb{R}^{d}$ attracted considerable attention with the high-dimensional variant reiterated recently by Alon and Kupavskii.
We prove that there exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{d}$ that have arbitrarily large girth. This improves upon a series of results due to Kupavskii; Sagdeev; and Sagdeev and Raigorodskii and gives the first bound in which the base of the exponent does not tend to one with the girth. In addition, our construction can be made explicit which allows us to answer in a strong form a question of Kupavskii.
Our arguments show graphs of large chromatic number and high girth exist in a number of other geometric settings including diameter graphs and orthogonality graphs.
1 Introduction
Chromatic number and girth are two of the perhaps most fundamental graph parameters, the study of which has attracted a vast amount of attention over the years. The chromatic number of a graph is the minimum number of colours needed to colour its vertices so that no two adjacent vertices receive the same colour (such a colouring is referred to as proper), and the girth of a graph is equal to the length of a shortest cycle in the graph. One of the most classical directions of study of these two parameters explores the relation between them. In particular, the question of whether one can obtain some amount of control of the chromatic number in terms of the girth has a very long and illustrious history. Indeed, at first glance, it is highly non-trivial to construct graphs with very high chromatic number, where this can not be witnessed locally. The local behaviour is captured quite well by the assumption of having large girth since this is equivalent to saying that the graph needs to locally look like a forest and any forest has chromatic number at most two.
That graphs with arbitrarily large girth and chromatic number do exist is a classical result of Erdős [12] from 1958, which features in essentially any course on the probabilistic method in combinatorics. Numerous alternative constructions with a variety of interesting further features were found over the years. Let us highlight the first explicit construction due to Lovász from 1966 [26], the extension to hypergraphs by Erdős and Hajnal [13] from the same year and the construction of Ramanujan graphs by Lubotzky, Phillips and Sarnak [27] from 1988, which besides high girth and chromatic number have a variety of other useful features associated with expander graphs and have had a tremendous impact in a variety of areas over the years. We point the reader to surveys by Nes̆etr̆il [30] and Raigorodskii and Shabanov [41] collecting just some of the very rich history surrounding this question.
The study of the chromatic number of a variety of geometric graphs is equally as classical a topic. For example, the Hadwiger-Nelson problem [19], dating back to 1945, asks for the maximum chromatic number of a unit distance graph in $\mathbb{R}^{d}$ while the classical Borsuk Problem [3], dating back to 1933, asks for the maximum chromatic number of a diameter graph in $\mathbb{R}^{d}$. Here a graph is said to be a unit distance graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct points in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to points at a unit distance. A graph is said to be a diameter graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct points in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to a pair of points at the maximum distance. We note that in both settings one may assume graphs involved are finite via the classical de-Bruijn-Erdős Compactness Theorem [6].
In the unit distance setting, in their remarkable 1981 paper essentially introducing the algebraic method in combinatorics, Frankl and Wilson [17] proved a conjecture of Erdős that there exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number exponential in $d$, improving upon a sequence of progressively better polynomial bounds (see e.g. the survey [44] for the full history). The Frankl-Wilson construction has been slightly improved since by Raigorodskii who holds the current record of $\chi(G)\geq(1.239+o(1))^{d}$ for a unit distance graph in $\mathbb{R}^{d}$.
In the diameter graph setting no bound beyond $\chi(G)=d+1$ obtained by the regular simplex was known. It was widely believed and is referred to as the Borsuk Conjecture that $\chi(G)\leq d+1$ for any diameter graph in $\mathbb{R}^{d}$, based in part on numerous results establishing this for special configurations of points. This was disproved in a stunning fashion by Kahn and Kalai in [21] who constructed diameter graphs with chromatic number growing exponentially in $\sqrt{d}$. There have been several improvements in the base, as well as a considerable effort in trying to prove any non-trivial bounds for small $d$, but constructing a diameter graph with exponential chromatic number remains a very interesting open problem. The current record is $\chi(G)\geq(1.225+o(1))^{\sqrt{d}}$ for $G$ a diameter graph in $\mathbb{R}^{d}$, due to Raigorodskii [42]. See the survey [43] for a more complete history and plenty of related problems.
A very natural next question, with the goal of understanding further what kind of substructures one can find in geometric graphs, is to ask whether one can find geometric graphs of high chromatic number and high girth. This question is quite classical in its own right and traces its origins to a problem posed by Erdős in 1975 [16] asking whether a unit distance graph in $\mathbb{R}^{2}$ with chromatic number $4$ and girth at least $4$ exists. Soon after he reiterated the problem saying it is likely that such a graph does not exist. Wormald [49] disproved this in 1978 by exhibiting such a graph on 6448 vertices. A construction on 23 vertices was found in 1996 by Hochberg and O’Donnell [20]. Erdős in the same 1975 paper also asked the more general question of whether unit distance graphs in $\mathbb{R}^{2}$ with chromatic number $4$ and arbitrarily large girth exist. This was only settled in 1999 by O’Donnell [33, 34], see also a recent survey of Graham [18]. Natural higher dimensional extensions were first considered by Raigorodskii [39] and Rubanov [45] in 2007. For more details and a plethora of related results see e.g. the survey by Raigorodskii [40].
Kupavskii [23] in 2012 proved that for each positive integer $k$ there exists some $c_{k}>1$ such that there are unit distance graphs in $\mathbb{R}^{d}$ with girth at least $k$ and chromatic number at least $(c_{k}+o(1))^{d}$. Alon and Kupavskii [2] reiterate the question of what can be said about the chromatic number of such graphs. Sagdeev [47] gave the first explicit estimate for $c_{k}$, namely that $c_{k}\geq 1+1/{k^{4+o(1)}}$. The dependency of $c_{k}$ on $k$ was subsequently improved to $c_{k}\geq 1+1/{k^{2+o(1)}}$ in [48] and by a further constant factor (implicit in $1/{k^{2+o(1)}}$) by Sagdeev and Raigorodskii in [46]. We prove a first universal exponential lower bound, without any dependency on $k$.
Theorem 1.
There exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{d}$ that have arbitrarily large girth.
Our bound even improves the best-known bounds for small values of $k$, e.g. in perhaps the most classical case of lower bounding the chromatic number of unit distance triangle-free graphs in $\mathbb{R}^{d}$ the previously best bound was $(1.058+o(1))^{d}$ [7]. The bounds on chromatic number of unit distance graphs in which we forbid a clique of size $k$ rather than all short cycles has also been considered, see [7, 25]. The bound in Theorem 1 improves the best-known lower bounds for the chromatic number of unit distance graphs in $\mathbb{R}^{d}$ with clique number at most $7$.
Quite a bit of focus over the years was afforded to finding explicit constructions of high girth and high chromatic number graphs, ever since the first such construction was found by Lovász in 1966 [26], in large part motivated by the idea that such graphs should exhibit somewhat random-like behaviour so could be useful in a variety of derandomisation arguments. Explicit constructions of graphs with large girth and chromatic number are now known in some geometric settings, including as box and line intersection graphs in $\mathbb{R}^{3}$ [4], as disjointedness graphs of short polygonal chains in the plane [35], and as tangency graphs of circles in the plane [5]. Our construction behind Theorem 1 is actually more flexible than previous, weaker ones in the unit distance graph setting and can be made explicit. This answers in quite a strong form a question of Kupavskii who asked to prove explicitly his result that unit distance graphs in $\mathbb{R}^{d}$ with girth at least $k$ and chromatic number at least $(c_{k}+o(1))^{d}$ exist for some $c_{k}>1$.
Our ideas can also be used to find graphs with high chromatic number and high girth in other geometric settings. The question of constructing high chromatic number and high girth diameter graphs in $\mathbb{R}^{d}$ for large $d$ has been raised recently by Prosanov [38], motivated by a result of Kupavskii and Polyanski [24]. We show that indeed strong counterexamples to Borsuk’s Conjecture can be found with arbitrarily large girth.
Theorem 2.
There exist diameter graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{\sqrt{d}}$ that have arbitrarily large girth.
Another setting, underlying and connecting the previous two results, is that of orthogonality graphs. Here a graph is said to be an orthogonality graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct vectors in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to orthogonal vectors.
Theorem 3.
There exist orthogonality graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{d}$ that have arbitrarily large girth.
In Section 2 we shall give a self-contained proof of Theorem 3 with a somewhat weaker bound for the base of the exponent. We also prove this weaker bound for Theorem 1 and Theorem 2 and more generally show that Theorem 3 implies Theorem 1 and Theorem 2. Then in Section 3 we give some more general black-box arguments that can be combined with state-of-the-art bounds to obtain Theorem 3 (and thus Theorem 1 and Theorem 2). In this section, we make our construction explicit by replacing a probabilistic sub-sampling step from Section 2 with an explicit construction for the desired sub-sampling (see Theorem 7). Lastly, in Section 4 we give some concluding remarks and discuss further open problems.
2 A self-contained argument
In this section, we will give a self-contained proof of Theorem 3 with a slightly weaker bound, in the interest of keeping the arguments as simple as possible. The goal is to showcase the wonderful interplay between algebraic and probabilistic ideas happening in the background. Since many of these ideas are by now standard in the area, experts may wish to skip to the following section where we extract the main ideas in the hope of making them more easily applicable in other settings and use them combined with existing state-of-the-art results to prove the bounds stated in the previous section.
Let $g$ be an arbitrary integer and let $p>2$ be a prime. We let $d=8p$ and will construct a unit distance graph in $\mathbb{R}^{d}$ with girth at least $g$ and chromatic number at least $\left(1.067+o(1)\right)^{d}$. [^1] We note that for $d$ not of this form, one could take $p$ to be the largest prime smaller than $d/8$ and use the known results on density of primes [11] to conclude the loss involved in using the construction in $8p$ dimensions only impacts the $o(1)$ term.
Let us now describe the construction. For a positive integer $m$, we denote the set $\{1,\ldots,m\}$ by $[m]$. In our construction, $m$ shall be some large integer to be chosen later. The vertex set of our graph will be
$V:=\left\{\left(\textbf{v}\cdot\cos\frac{2\pi j}{m},\textbf{v}\cdot\sin\frac{2\pi j}{m}\right)\mid\textbf{v}\in\{\pm 1\}^{d/2},\textbf{v}_{d/2}=1,\sum_{i=1}^{d/2}\textbf{v}_{i}=0,j\in[m]\right\}\subseteq\mathbb{R}^{d}.$
The edge set will be obtained by randomly subsampling orthogonal pairs of vectors in $V$. Let us however first introduce some notation. We let $V^{\prime}:=\{\textbf{v}\mid\textbf{v}\in\{\pm 1\}^{d/2},\textbf{v}_{d/2}=1,\sum_{i=1}^{d/2}\textbf{v}_{i}=0\}$ and for $\textbf{v}\in V^{\prime},j\in[m]$ we let $\textbf{v}(j)=\left(\textbf{v}\cdot\cos\frac{2\pi j}{m},\textbf{v}\cdot\sin\frac{2\pi j}{m}\right)$. So that $V=\{\textbf{v}(j)\mid\textbf{v}\in V^{\prime},j\in[m]\}$. One can think of $V$ as consisting of $m$ spaced out rotations of each $\textbf{v}\in V^{\prime}$ in the plane defined by $(\textbf{v},\textbf{0})$ and $(\textbf{0},\textbf{v})$ in $\mathbb{R}^{d}$.
The key property of $V$ is that if v is orthogonal to u for $\textbf{v},\textbf{u}\in V^{\prime}$ then $\textbf{v}(i)$ is orthogonal to $\textbf{u}(j)$ for any $i,j\in[m]$. So if one denotes by $G^{\prime}$ the graph with vertex set $V^{\prime}$ and edge set being orthogonal pairs in $V^{\prime}$, then the blow-up graph $G^{\prime m}$ (which is the graph obtained from $G^{\prime}$ by replacing each vertex with an independent set of size $m$ and each edge with a $K_{m,m}$ subgraph between the two corresponding independent sets of size $m$) is an orthogonality graph in $\mathbb{R}^{d}$ with vertex set $V$. Armed with this observation our plan is to show that $\chi(G^{\prime})$ is exponential in $d$ and that thanks to that if we subsample the edges of $G^{\prime m}$ we can preserve the same chromatic number while removing all cycles of length shorter than $g$.
That the first property holds follows from a slightly modified argument of Nilli [32], which is, in turn, a slight modification of the one used by Frankl and Wilson [17] when first establishing exponential growth of the chromatic number of the unit distance graph of $\mathbb{R}^{d}$.
Theorem 4.
$\chi(G^{\prime})\geq\left({27}/{16}+o(1)\right)^{p}\geq\left(1.067+o(1)\right)^{d}.$
  • Proof.
[…]
    Finally, as in any proper colouring, every colour class spans no edges, this implies that we need at least
    $\frac{|V^{\prime}|}{2\binom{4p}{p-1}}=\frac{\binom{4p}{2p}}{4\binom{4p}{p-1}}\geq\left({27}/{16}+o(1)\right)^{p}\geq\left(1.067+o(1)\right)^{d}$
    colours to properly colour $G^{\prime}$, as claimed. ∎
We now turn to the second part of the argument, namely increasing the girth. Let $k=\chi(G^{\prime})$ and let $m=2^{8pg}\cdot(4k)^{4g}\cdot(k-1)^{2}$. As we already argued $G^{\prime m}$ is an orthogonality graph in $\mathbb{R}^{d}$. Let $n=m\cdot 2^{4p}\geq|V(G^{\prime m})|.$ Let $G_{0}\sim G^{\prime m}(q)$ be the random graph obtained by including every edge of $G^{\prime m}$ in $G_{0}$ with probability $q:=1/n^{1-1/(2g)}$, independently from all other edges. The expected number of cycles of length $\ell$ in $G_{0}$ is at most $\binom{n}{\ell}\cdot\frac{\ell!}{2\ell}\cdot q^{\ell}\leq(nq)^{\ell}/2=n^{\ell/(2g)}/2$. So, the expected number of cycles of length at most $g$ is at most $\sum_{\ell=3}^{g}n^{\ell/(2g)}/2\leq n^{1/2}.$ By Markov’s inequality, we can conclude that the probability there are more than $2\sqrt{n}$ cycles in $G_{0}$ is less than $1/2$.
On the other hand, if we write $m^{\prime}:=\frac{m}{k-1}$ we want to show that with probability less than $1/2$ there exist adjacent $v,u\in V(G^{\prime})$ and a pair of subsets $V,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively with at most $2\sqrt{n}$ edges between $V$ and $U$. To see that note that there are at most $e(G^{\prime})\cdot\binom{m}{m^{\prime}}^{2}$ choices for $V$ and $U$ and the chance there are at most $2\sqrt{n}$ edges between them is at most $\binom{m^{\prime 2}}{2\sqrt{n}}(1-q)^{m^{\prime 2}-2\sqrt{n}}$. So the desired statement holds by a union bound as
$e(G^{\prime})\cdot\binom{m}{m^{\prime}}^{2}\cdot\binom{m^{\prime 2}}{2\sqrt{n}}(1-q)^{m^{\prime 2}-2\sqrt{n}}\leq 2^{8p}\cdot 2^{2m}\cdot 2^{4\sqrt{n}\log m^{\prime}}\cdot 2^{-qm^{\prime 2}+1}<\frac{1}{2},$
where the final inequality follows since $qm^{\prime 2}\geq m^{1+1/(2g)}/(2^{4p}k^{2})\geq 4m\geq 8p+2m+4\sqrt{n}\log m+2.$
So there exists an outcome in which graph $G_{0}$ has at most $2\sqrt{n}$ cycles of length up to $g$ and for any adjacent $v,u\in V(G^{\prime})$ and for any pair of subsets $V,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively, there are more than $2\sqrt{n}$ edges between $V$ and $U$. We obtain our graph $G$ by removing an arbitrary edge from every cycle of length up to $g$ in $G_{0}$ to obtain a graph with girth at least $g$ and with the property that there is an edge between any pair of subsets $V,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively.
Suppose that $\chi(G)<\chi(G^{\prime})=k$. This means there exists a colouring $c:V(G)\to[k-1]$ such that there are no adjacent vertices with the same colour. Now colour every vertex $v$ of $G^{\prime}$ by the majority colour according to $c$ appearing on the $m$ vertices corresponding to $v$ in $V(G^{\prime m})=V(G).$ This produces a $(k-1)$-colouring of $G^{\prime}$ which can not be proper, as $\chi(G^{\prime})=k$, and hence there must exist adjacent $v,u\in V(G^{\prime})$ given the same colour. By definition of the colouring, there exist two subsets $V,U$ of $\frac{m}{k-1}=m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively, all given the same colour in $c$. But there exists an edge between these $V$ and $U$ by our construction, which contradicts the definition of $c$.
This completes the proof of Theorem 3 with a somewhat weaker base of the exponential (being 1.067). Next, we obtain Theorem 1 and Theorem 2 with the same somewhat weaker base of the exponentials. The same arguments also show that Theorem 3 implies Theorem 1 and Theorem 2.
Since all the vertices in $V$ have the same norm $||\textbf{v}(j)||^{2}=||\textbf{v}||^{2}\left(\cos^{2}\frac{2\pi j}{m}+\sin^{2}\frac{2\pi j}{m}\right)=||\textbf{v}||^{2}=4p$ we know that any orthogonal pair $\textbf{v}(i),\textbf{v}(j)$ satisfy $||\textbf{v}(i)-\textbf{v}(j)||^{2}=8p$. So after scaling all the vectors by a factor of $\sqrt{8p}$ we obtain our graph $G$ is also a unit distance graph, proving Theorem 1 with a smaller base of the exponential.
[…]
3 Black-box arguments
In this section, we extract the main ideas used in the previous one and apply them together with state-of-the-art lower bounds on the chromatic number without any girth restrictions to prove our results.
The first one is that by doubling the number of dimensions one can embed an arbitrary blow-up of an orthogonality graph. Recall that for a graph $G$ and a positive integer $m$, the blow-up graph $G^{m}$ is the graph obtained from $G$ by replacing each vertex with an independent set of size $m$ and each edge with a $K_{m,m}$ subgraph between the two corresponding independent sets of size $m$.
Proposition 5.
For any $m\in\mathbb{N}$ and orthogonality graph $G$ in $\mathbb{R}^{d}$, $G^{m}$ is an orthogonality graph in $\mathbb{R}^{2d}$.
  • Proof.
    Let $G$ be an orthogonality graph in $\mathbb{R}^{d}$ with representation $V\subseteq\mathbb{R}^{d}$. For each $\textbf{v}\in V$, let $\textbf{v}^{\prime}=(\textbf{v},\textbf{0})\in(\mathbb{R}^{d})^{2}=\mathbb{R}^{2d}$, and $\textbf{v}^{\prime\prime}=(\textbf{0},\textbf{v})\in(\mathbb{R}^{d})^{2}=\mathbb{R}^{2d}$. Then, for each orthogonal pair $\textbf{u},\textbf{v}\in V$, we have that $\operatorname{span}_{\mathbb{R}}\{\textbf{u}^{\prime},\textbf{u}^{\prime\prime}\}$ and $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ are orthogonal planes in $\mathbb{R}^{2d}$. For each $\textbf{v}\in V$, choose distinct points $\textbf{v}_{1},\ldots,\textbf{v}_{m}$ in $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ so that the points are also distinct between different $\textbf{v}\in V$. Now observe that for every orthogonal pair $\textbf{u},\textbf{v}\in V$, since $\operatorname{span}_{\mathbb{R}}\{\textbf{u}^{\prime},\textbf{u}^{\prime\prime}\}$ and $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ are orthogonal planes, we have for every $1\leq i,j\leq m$ that $\textbf{u}_{i}$ is orthogonal to $\textbf{v}_{j}$. Therefore, $G^{m}$ is an orthogonality graph in $\mathbb{R}^{2d}$. ∎
The probabilistic part of the argument can be captured by the following theorem see e.g. [50]. An alternative probabilistic proof, based on the Lovász Local Lemma, of a somewhat stronger result can be found in [28].
Theorem 6 (Zhu [50]).
For every graph $G$ and positive integer $g$, there exists a positive integer $m$ such that $G^{m}$ contains a subgraph $G^{\prime}$ with girth at least $g$ and $\chi(G^{\prime})=\chi(G)$.
In the following theorem, we prove a slight strengthening which comes with an explicit proof. Loosely speaking our result says that given any graph $G$ we can find a subgraph $G^{\prime}$ of some blow-up graph of $G$ with large girth and with the property that any proper $k$-colouring of $G^{\prime}$ gives rise to a proper $k$-coloring of $G$ (with the property that a colour of $v\in V(G)$ appears as the colour of some vertex in the part corresponding to $v$ in the blow-up).
Our argument is a modification of one of Nešetřil and Rödl [29] that they used to construct graphs with large girth and chromatic number, which itself is a modification of one of Tutte [8, 9] who constructed triangle-free graphs with large chromatic number. As part of our argument, we will need the existence of (explicit) hypergraphs with high girth and chromatic number. A cycle of length $\ell\geq 2$ in a hypergraph $H$ is a tuple $(F_{1},\ldots,F_{\ell})$ of distinct hyperedges such that there exist distinct elements $v_{1},\ldots,v_{\ell}\in V(H)$ with $v_{i}\in F_{i}\cap F_{i+1}$ for $11\leq i\leq\ell-$ 1, and $v_{\ell}\in F_{\ell}\cap F_{1}$. The girth of a hypergraph is equal to the length of a shortest cycle. The chromatic number of a hypergraph is defined as the minimum number of independent sets (sets not containing any edge fully) needed to partition its vertex set. The existence of hypergraphs of arbitrary uniformity with high girth and chromatic number follows from an immediate strengthening of the classical probabilistic argument of Erdős [12] and is due to Erdős and Hajnal [13]. Multiple explicit constructions have been found since see e.g. [1, 10, 14, 22, 26, 31, 36, 37]. Given a vertex $u$ of a graph $G$, we let $N_{G}(u)$ denote the neighbourhood of $u$ in $G$.
Theorem 7.
For every graph $G$ and positive integer $g$ there exists a graph $G^{\prime}$ with girth at least $g$ and a surjective homomorpishm $h$ from $G^{\prime}$ to $G$ such that for any proper $k$-colouring $c^{\prime}$ of $G^{\prime}$, there exists a proper $k$-colouring $c$ of $G$ such that $c(v)\in c^{\prime}(h^{-1}(v))$ for all $v\in V(G)$. In particular, $\chi(G^{\prime})=\chi(G)$.
  • Proof.
    If $|V(G)|\leq 2$ the theorem is immediate since we may just take $G^{\prime}=G$ and let $h$ to be the identity. So, with this observation serving as a basis, we shall argue inductively. Let $u$ be a vertex of $G$, then we may assume that the theorem holds for $G-u$. Now, let $G_{u}$ be a graph with girth at least $g$ and $h_{u}$ a surjective homomorpishm from $G_{u}$ to $G-u$ such that for any proper $k$-colouring $c_{u}$ of $G_{u}$, there exists a proper $k$-colouring $c$ of $G-u$ such that $c(v)\in c_{u}(h_{u}^{-1}(v))$ for all $v\in V(G-u)$.
    Let $n=|h_{u}^{-1}(N_{G}(u))|$, and let $H$ be an auxiliary $n$-uniform hypergraph with chromatic number at least $k+1$ and girth at least $g$. For each $F\in E(H)$, let $G_{F}$, $h_{F}$ be copies of $G_{u},h_{u}$, where the $G_{F}$’s are taken to be vertex disjoint.
    We construct our desired $G^{\prime}$ by taking a vertex disjoint union of all $G_{F}$ together with a disjoint independent set $U$ whose vertices we identify with $V(H)$. In addition, for each $F\in E(H)$, we add an arbitrary matching between the vertices of $U$ corresponding to those of $F$ and $h_{F}^{-1}(N_{G_{F}}(u))\subseteq V(G_{F})$. We define our desired $h:V(G^{\prime})\to V(G)$ by letting $h(v)=h_{F}(v)$ for all $v\in V(G_{F})$ and $h(v)=u$ for all $v\in U$. Since $h_{F}$ is a surjective homomorphism to $G-u$, so is $h$ to $G$.
    Consider a cycle $C$ in $G^{\prime}$. If $C$ is contained in some $G_{F}$, then clearly $C$ has length at least $g$. Otherwise, if $C$ is not contained in some $G_{F}$, then $C$ must contain some vertices $u_{1},\ldots,u_{r}$ of $U$ which we label according to the order in which they appear on $C$ (choosing $u_{1}$ and direction arbitrarily). Observe that for each $1\leq r\leq i$, the internal vertices of the subpath of $C$ between $u_{i}$ and $u_{i+1}$ (taking $u_{r+1}=u_{1}$) are contained in a unique $G_{F_{i}}$, where $F_{i}\in E(H)$, with $u_{i},u_{i+1}\in F_{i}$. This follows since $u_{i}$ and $u_{i+1}$ can not be adjacent on the cycle (since they belong to $U$ which is independent) and since there are no edges between distinct $G_{F}$. Furthermore, for each $i$ we have that $F_{i}\not=F_{i+1}$ (again taking $F_{r+1}=F_{1}$), since the edge from $u_{i+1}$ to $G_{F_{i}}$ is unique and can be traversed only once by $C$. Since $H$ has girth at least $g$, it follows that $r\geq g$. Therefore, in either case, $C$ has length at least $g$. So, $G^{\prime}$ has girth at least $g$.
    Consider some proper $k$-colouring $c^{\prime}$ of $G^{\prime}$. Restriction of $c^{\prime}$ to $U$ gives a $k$-coloring of $H$, which has chromatic number at least $k+1$, so some edge of $H$ must be monochromatic in this coloring. Namely, there is some colour $a$ and some $F\in E(H)$ such that $c^{\prime}(F)=\{a\}$. Therefore, since $c^{\prime}$ was proper and by our construction, there is a matching between $F\subseteq U$ and $h_{F}^{-1}(N_{G_{F}}(u))$ we know $c^{\prime}(h_{F}^{-1}(N_{G_{F}}(u)))\subseteq\{1,\ldots,k\}\backslash\{a\}$. So, by applying the inductive assumption to the restriction of $c^{\prime}$ to $G_{F}$ we obtain a proper $k$-colouring $c$ of $G-u$ such that $c(v)\in c^{\prime}(h^{-1}(v))$ for all $v\in V(G-u)$. Since $a\notin c^{\prime}(h_{F}^{-1}(N_{G_{F}}(u)))$ we can extend $c$ by letting $c(u)=a$ to obtain a desired proper $k$-coloring of $G$.
    This in particular implies that if there exists a proper $k$-colouring of $G^{\prime}$ there also exists a proper $k$-coloring of $G$ so $\chi(G^{\prime})\geq\chi(G)$. On the other hand, since there is a homomorphism from $G^{\prime}$ to $G$, we have that $\chi(G^{\prime})\leq\chi(G)$, so $\chi(G^{\prime})=\chi(G)$. ∎
Remark. In private communication Alon suggested an alternative way to prove Theorem 6 explicitly. The basic idea is to start with a high-girth Ramanujan graph (such as an LPS-graph [27]) on $|G|m$ vertices. One can then assign to each vertex of $G$ a disjoint set of $m$ of its vertices and remove any edges within parts or between parts corresponding to non-adjacent vertices of $G$. One can then show using a similar argument as we did with random subsampling that the chromatic number of this graph will remain the same as that of $G$, provided $m$ is chosen to be sufficiently large.
The final ingredient is the state-of-the-art construction of orthogonality graphs with large chromatic number due to Raigorodskii [42]. We note that the construction is similar in spirit to the one we presented in the previous section with the key additional idea being to use vectors with coordinates in $\{-1,0,1\}$ instead of just $\{-1,1\}$.
Theorem 8 (Raigorodskii [42]).
There exist orthogonality graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.2255+o(1))^{d}$.
By combining Theorem 8 with Proposition 5 and Theorem 7, we obtain Theorem 3 (since $\sqrt{1.2255}=1.107...$). As discussed at the end of Section 2, Theorem 3 then implies Theorem 1 and 2.
4 Concluding remarks
In this paper, we proved that there exist, unit distance and orthogonality graphs in $\mathbb{R}^{d}$ and diameter graphs in $\mathbb{R}^{d^{2}}$ with chromatic number exponential in $d$ and with arbitrarily large girth.
We note that there is another natural notion of unit distance graphs (and similarly for diameter and orthogonality graphs), sometimes referred to as faithful or induced unit distance graphs in which one insists that any pair of points in $\mathbb{R}^{d}$ at unit distance do make an edge. Unfortunately, due to the fact we subsample edges, our arguments do not apply in this setting and it remains a very interesting open problem to find faithful unit distance graphs with large girth and chromatic number. We point the reader to a paper of Alon and Kupavskii [2] collecting the differences between the two notions.
We note also that our Proposition 5 allows us to embed (faithfully) in $\mathbb{R}^{2d}$ a blow-up of any (faithful) orthogonality graph in $\mathbb{R}^{d}$. It could be interesting to determine how much we need to increase the dimension, if at all, for this to hold. One does need to increase the number of dimensions when $d\leq 3$ by the results on the classical Erdős Unit Distance Problem [15] but for $d\geq 4$ it might be possible to embed the blow-up in the same space. If so this would slightly improve the bases of exponentials in our results.
Another interesting direction which arises is the fact that not all unit distance graphs are orthogonality graphs (of the same dimension), so it might be interesting to determine whether it is possible to prove some analogue of Proposition 5 for unit distance graphs. If so then one could use the slightly better constructions known for unit distance graphs compared to orthogonality ones and obtain a very slight improvement over our Theorem 1. Understanding both this and the previous question might help in improving the bounds on the Borsuk Problem without any girth conditions.
Acknowledgements
We thank Noga Alon for sharing his alternative explicit proof of Theorem 6 and thank Jim Geelen for helpful discussions on topics related to Theorem 6.